OpenStudy (anonymous):

Find the derivative for y^4 + x^3y^3 = 1 + e^2x y^3

OpenStudy (anonymous):

is implicit: D(y^4) + D(x^3y^3) = D(1) + D(e^2xY^3) 4y^3 + 3x^2y^3+3y^2X^3 = 2e^2xy^3+e^2x3y^2 group all the y equations on the left side and the x equations on the right then; divide to get derivative of y.